# Cross-View UAV Geo-Localization with Precision-Focused Efficient Design: A Hierarchical Distillation Approach with Multi-view Refinement

### v1

@@ Abstract
Cross-view geo-localization (CVGL) enables UAV localization by matching aerial images to geo-tagged satellite databases, which is critical for autonomous navigation in GNSS-denied environments. However, existing methods rely on resource-intensive fine-grained feature extraction and alignment, where multiple branches and modules significantly increase inference costs, limiting their deployment on edge devices. We propose Precision-Focused Efficient Design (PFED), a resource-efficient framework combining hierarchical knowledge transfer and multi-view representation refinement. This innovative method comprises two key components: 1) During training, Hierarchical Distillation paradigm for fast and accurate CVGL (HD-CVGL), coupled with Uncertainty-Aware Prediction Alignment (UAPA) to distill essential information and mitigate the data imbalance without incurring additional inference overhead. 2) During inference, an efficient Multi-view Refinement Module (MRM) leverages mutual information to filter redundant samples and effectively utilize the multi-view data. Extensive experiments show that PFED achieves state-of-the-art performance in both accuracy and efficiency, reaching 97.15% Recall@1 on University-1652 while being over more efficient in FLOPs and faster than previous top methods. Furthermore, PFED runs at 251.5 FPS on the AGX Orin edge device, demonstrating its practical viability for real-time UAV applications. The project is available at https://github.com/SkyEyeLoc/PFED.
@@ I Introduction
To improve the cross-view matching precision, the field has rapidly evolved from early methods using handcrafted descriptors to dominant deep learning paradigms built on Siamese or Triplet networks [1, 2, 3, 4, 5]. More recently, Vision Transformers (ViTs) [6] and their variants, such as TransGeo [7] and FSRA [8], have set high performance standards by leveraging self-attention to learn powerful, view-invariant global representations. However, despite these advancements, several critical challenges remain that hinder the deployment of these models in practical, real-world scenarios.
Firstly, the pursuit of higher accuracy has led to an escalating computational burden. Many state-of-the-art methods achieve superior performance by incorporating complex auxiliary branches, resource-demanding cross-view alignment strategies. As shown in Figure 1, while models like CCR [9] and Mean [4] achieve high recall, they do so with immense computational overhead (e.g., over 90G FLOPs). This paradigm is fundamentally ill-suited for resource-constrained edge platforms like drones, where memory, power, and latency are primary concerns, creating a significant gap between algorithmic advancement and practical applicability.
@@
Based on the above analysis, a straightforward idea is to simply deploy a lightweight network, but this approach typically leads to a significant performance collapse. To bridge this performance gap, we propose the Precision-Focused Efficient Design (PFED), a novel framework that fundamentally redefines the accuracy-efficiency trade-off by shifting complexity from the inference stage to the training stage. As illustrated in Figure 2, our approach comprises two key innovations. 1) During training, we introduce Hierarchical Distillation for Cross-View Geo-Localization (HD-CVGL). This paradigm first employs Fine-Grained Inverse Self-distillation (FISD) to compel the model’s final, semantically rich layer to retain the critical spatial details discovered by its shallower layers. Concurrently, our Uncertainty-Aware Prediction Alignment (UAPA) mechanism adaptively balances the learning between drone and satellite branches based on predictive uncertainty, fostering a more stable and robust training process. Finally, we perform cross-distillation from a domain-specialized foundation model to distill generalized knowledge into our lightweight student network, boosting its precision without incurring any additional inference overhead. 2) During inference, we propose a lightweight Multi-view Refinement Module (MRM). This post-processing method leverages mutual information to intelligently select a small, diverse, and highly informative subset of drone views. By aggregating the features from only these key views, MRM significantly enhances localization accuracy while drastically reducing the feature matching cost, enabling efficient and precise multi-view utilization.
@@
(1) We propose the Precision-Focused Efficient Design (PFED) paradigm, which resolves the accuracy-efficiency dilemma in CVGL by concentrating model complexity in the training phase, resulting in a highly accurate yet computationally lightweight inference model.
@@
(3) We propose the Multi-view Refinement Module (MRM), an efficient post-processing method, and theoretically demonstrate based on manual information how it enhances localization by optimally selecting and fusing multi-view drone information while minimizing feature matching overhead.
(4) The empirical results on the University-1652 and SUES-200 benchmarks demonstrate that PFED establishes a new state-of-the-art in both accuracy and efficiency. Additional deployment tests on edge devices further validate PFED’s real-time performance capabilities.
@@ II Related work
II-A Cross-view UAV Geo-localization
Cross-View Geo-Localization involves matching a query image against a geo-referenced database from another viewpoint, a foundational challenge within the field of computer vision. Early methods relied on handcrafted local descriptors like SIFT [15] or global scene descriptors like HoG [16], but showed limited robustness to the extreme cross-view domain gap. The paradigm shifted with deep metric learning, where Siamese or Triplet networks [17, 18] learn a shared, view-invariant embedding space. To explicitly address geometric inconsistencies, some works integrate geometric priors, notably using a polar transform to align satellite and ground-level perspectives [19, 20] before feature comparison. While ground-to-satellite matching [21, 22] sets the foundation, the unique oblique perspective of Unmanned Aerial Vehicles (UAVs) introduces a more severe and spatially non-uniform multi-view set of geometric challenges.
UAV-to-Satellite Geo-Localization. This task is particularly challenging due to the multi-view oblique and low-altitude perspective of UAVs, which creates significant domain gaps between platforms. Following the establishment of key benchmarks [1, 12], research has progressed from improving feature robustness with attention mechanisms [23] to the now-dominant Vision Transformer (ViT) architectures. ViTs like TransGeo [7] and FSRA [24] leverage self-attention to learn powerful global representations, setting high performance standards. More recently, the field has explored more efficient and powerful backbone architectures. For instance, several works have successfully employed ConvNeXt [25] to extract highly discriminative global features, achieving a strong balance of performance and efficiency [26]. This trend highlights a focus on backbone potency for feature extraction. However, most of contemporary methods [4, 27, 13] achieve higher accuracy by incorporating sophisticated auxiliary branches or modules, this paradigm is ill-suited for resource-constrained platforms. We follow a Precision-Focused Efficient Design (PFED) paradigm to empower a simple backbone. By strengthening its feature extraction capabilities without adding any inference-time modules, our model achieves competitive precision while drastically reducing the computational burden, making it ideal for deployment on platforms like UAVs.
@@ II-B Multi-view Refinement
Traditional UAV-based geo-localization relies on direct matching between individual query images and a reference database [28, 3, 8]. However, these approaches struggle with significant viewpoint discrepancies caused by factors such as occlusions from structures or vegetation, and the diverse perspectives captured by UAVs operating at varying altitudes and angles, making accurate location recognition increasingly challenging. Recent works have recognized the importance of leveraging multiple drone views to improve localization accuracy with approaches using 3D reconstruction [29, 30, 31] to represent scenes from multi-view observations and iteratively refining camera poses to align rendered views with satellite imagery. This multi-view fusion paradigm [32] demonstrates significant improvements in localization performance by exploiting the rich geometric and appearance information contained across different viewpoints.
However, despite their theoretical advantages, these multi-view approaches impose substantial computational burdens that severely limit their practical UAV onboard applicability. In contrast to these complex approaches with substantial computational barriers, we propose an efficient multi-view refinement module (MRM) that operates as a lightweight post-processing method. Rather than constructing expensive 3D representations or performing iterative rendering optimization, our method directly aggregates selected features from multiple drone viewpoints through Mutual Information (MI).
@@
Mutual information has emerged as a principled criterion for selecting informative views in a wide variety of complex systems. In 3D reconstruction, mutual information has guided next-best-view selection effectively. Recent work proposes gaussian splatting shannon mutual information for real-time assessment of visual mutual information from novel viewpoints, enabling principled next-best-view selection without prior information [33]. For contrastive learning, Tian et al. [34] argue that we should reduce the mutual information between views while keeping task-relevant information intact, proposing the InfoMin principle. Similarly, in multi-view clustering, recent work [35] minimizes MI between common and view-specific representations to exploit inter-view complementary information while maximizing MI between refined and original representations to preserve principal information.
Existing UAV-based geo-localization methods either process all available views through expensive reconstruction-based methods [14, 29], missing the opportunity for principled information-theoretic view selection. In this work, we bridge this gap by proposing an efficient mutual information-based post-processing method that jointly optimizes view selection and feature aggregation for UAV-based geo-localization.
@@ III Proposed Method
In this section, we provide a comprehensive introduction to our proposed PFED. As illustrated in Figure 2, we follow a precision-focused efficient design in two stages: training and inference post-processing. During the training stage, we propose Hierarchical Distillation for CVGL (HD-CVGL). Initially, we employ Fine-Grained Inverse Self-distillation (FISD) based on the internal stages of an efficient student model. Concurrently, to address the challenge of data imbalance, we introduce Uncertainty-Aware Prediction Alignment (UAPA), achieving a more robust and stable train. Subsequently, guided by a fine-tuned foundation model, we perform cross-distillation to obtain a student model with higher precision. In the inference stage, we propose a Multi-view Refinement Module (MRM) to better leverage multi-view information and further reduce the time cost of feature matching.
@@ III-A Hierarchical Distillation for CVGL
The second step of our hierarchical framework is Cross-Distillation Training (CDT), a process designed to transfer knowledge from a large foundation model (DINOv2-base [40]) teacher to a lightweight ConvNeXt-tiny [25] student, as illustrated in Figure 2 (a). This approach adheres to the PFED paradigm by directly training the final student architecture without requiring any inference-time modifications. Critically, this teacher is not used off-the-shelf; it is first specialized for the target task through a parameter-efficient fine-tuning process on the University-1652 dataset, where only the final two Transformer blocks were made trainable. This approach creates an expert teacher that retains vast, general visual knowledge while acquiring high-level semantic understanding specific to our domain.
@@
III-B Multi-view Refinement Module (MRM)
@@
To address this challenge, as shown in Figure 4, we propose the Multi-view Refinement Module (MRM), a post-processing technique that intelligently selects an optimal subset with . The key insight is that not all views contribute equally to geo-localization accuracy: some perspectives capture more distinctive features, while others may be less informative due to occlusions or viewing angles. We formulate this view selection problem within an information-theoretic framework, maximizing the mutual information between landmark identities while ensuring spatial diversity.
@@
Information-Theoretic View Importance Assessment. Given extracted multi-view features for a landmark, the MRM quantifies each view’s information content through three complementary measures grounded in information theory. First, we approximate the marginal entropy of view features through log-variance:
@@
Computational Efficiency: By reducing the feature matching complexity from to , our MRM achieves a significant speedup. is the number of drone views of one landmark, this corresponds to a reduction in computational cost, rendering the approach viable for real-time applications.
@@ IV-A Implementation Details
Evaluation Protocol. Following standard practice in cross-view geo-localization [1, 7], we evaluate both retrieval accuracy and computational efficiency across all models
@@
Efficiency Metrics. We report three efficiency metrics: (1) Parameters: the total number of learnable parameters in millions; (2) FLOPs: floating-point operations computed for a single forward pass; (3) Processing Speed (FPS): in the feature extraction context, the number of UAV and satellite images processed per second. Note that we set the batch size to 16 to accommodate the multi-view UAV data processing.
@@
| LPN [2] [TCSVT’2021] | 62.39 M | 65.39 G | 218 | 77.71 | 80.80 | 90.30 | 78.78 |
| FSRA [8] [TCSVT’2021] | 53.16 M | 98.05 G | 100 | 85.50 | 87.53 | 89.73 | 84.94 |
| MCCG [28] [TCSVT’2023] | 56.65 M | 51.04 G | 313 | 89.40 | 91.07 | 95.01 | 89.93 |
| MuSe-Net [46] [PR’2024] | 82.90 M | 42.37 G | 405 | 74.48 | 77.83 | 88.02 | 75.10 |
| SCPNet [47] [TCSVT’2025] | - | - | - | 79.96 | 83.04 | 87.33 | 79.87 |
| TriSA [3] [TCSVT’2024] | 51.13 M | 43.18 G | 275 | 90.08 | 91.56 | 96.01 | 90.12 |
| Safe-Net [48] [TIP’2024] | 52.67 M | 24.58 G | 282 | 86.98 | 88.85 | 91.22 | 86.06 |
| SDPL [10] [TCSVT’2024] | 42.56 M | 69.71 G | 519 | 90.16 | 91.64 | 93.58 | 89.45 |
| SRLN [49] [TGRS’2024] | 193.03 M | - | - | 92.70 | 93.77 | 95.14 | 91.97 |
| Sample4Geo [13] [ICCV’2023] | 87.57 M | 90.24 G | 144 | 92.65 | 93.81 | 95.14 | 91.39 |
| CCR [9] [TCSVT’2024] | 156.57 M | 160.61 G | - | 92.54 | 93.78 | 95.15 | 91.80 |
| DAC [27] [TCSVT’2024] | 96.50 M | 90.24 G | 128 | 94.67 | 95.50 | 96.43 | 93.79 |
| MEAN [4] [TGRS’2025] | 36.50 M | 26.18 G | 307 | 93.55 | 94.53 | 96.01 | 92.08 |
| PFED (Ours) | 28.57 M | 4.45 G | 1022 | 93.87 | 94.83 | 95.72 | 92.57 |
| PFED (Ours w/ Post-process) | 28.57 M | 4.45 G | 1022 | 97.15 | 97.50 | 95.58 | 96.27 |

### v2

# MobileGeo: Exploring Hierarchical Knowledge Distillation for Resource-Efficient Cross-view Drone Geo-Localization

@@ Abstract
Cross-view geo-localization (CVGL) enables drone localization by matching aerial images to geo-tagged satellite databases, which is critical for autonomous navigation in GNSS-denied environments. However, existing methods rely on resource-intensive feature alignment and multi-branch architectures, incurring high inference costs that limit their deployment on mobile edge devices. We propose MobileGeo, a mobile-friendly framework designed for efficient on-device CVGL. MobileGeo achieves its efficiency through two key components: 1) During training, a Hierarchical Distillation (HD-CVGL) paradigm, coupled with Uncertainty-Aware Prediction Alignment (UAPA), distills essential information into a compact model without incurring inference overhead. 2) During inference, an efficient Multi-view Selection Refinement Module (MSRM) leverages mutual information to filter redundant views and reduce computational load. Extensive experiments demonstrate that MobileGeo outperforms previous state-of-the-art methods, achieving a 4.19% improvement in AP on University-1652 dataset while being over 5 more efficient in FLOPs and 3 faster. Crucially, MobileGeo runs at 251.5 FPS on an NVIDIA AGX Orin edge device, demonstrating its practical viability for real-time on-device drone geo-localization. The code is available at https://github.com/SkyEyeLoc/MobileGeo.
@@ I Introduction
To improve the cross-view matching precision, the field has rapidly evolved from early methods using handcrafted descriptors to dominant deep learning paradigms built on Siamese or Triplet networks [1, 2, 3, 4, 5]. More recently, Vision Transformers (ViTs) [6] and their variants, such as TransGeo [7] and FSRA [8], have set high performance standards by leveraging self-attention to learn powerful, view-invariant global representations. However, despite these advancements, several critical challenges remain that hinder the deployment of these models in practical, real-world mobile scenarios.
Firstly, the pursuit of higher accuracy has led to an escalating computational and resource burden. In Figure 2, many state-of-the-art methods achieve superior performance by incorporating complex auxiliary branches, resource-demanding cross-view alignment strategies. As shown in Figure 1, while models like CCR [11] and Mean [4] achieve high recall, they do so with immense computational overhead (e.g., over 90G FLOPs). This paradigm is fundamentally ill-suited for resource-constrained edge platforms, such as drones, where memory, power, and latency are critical concerns, creating a significant gap between algorithmic advancements and their practical applicability in mobile computing.
@@
Based on the above analysis, a straightforward idea is to simply deploy a lightweight network, but this approach typically leads to a significant performance collapse. To bridge this performance gap, we propose the MobileGeo, a novel mobile-friendly framework that fundamentally redefines the accuracy-efficiency trade-off by shifting complexity from the inference stage to the training stage. As illustrated in Figure 3, our approach comprises two key innovations. 1) During training, we introduce Hierarchical Distillation for Cross-View Geo-Localization (HD-CVGL). This paradigm first employs Fine-Grained Inverse Self-Distillation (FISD) to compel the model’s final, semantically rich layer to retain the critical spatial details discovered by its shallower layers. Concurrently, our Uncertainty-Aware Prediction Alignment (UAPA) mechanism adaptively balances the learning between drone and satellite branches based on predictive uncertainty, fostering a more stable and robust training process. Finally, we perform cross-distillation from a domain-specialized foundation model to distill generalized knowledge into our lightweight student network, boosting its precision without incurring any additional inference overhead. 2) During inference, we propose a lightweight Multi-view Selection Refinement Module (MSRM). This post-processing method leverages mutual information to intelligently select a small, diverse, and highly informative subset of drone views. By aggregating the key features, MSRM significantly enhances accuracy while drastically reducing the feature matching cost, enabling mobile-efficient multi-view utilization.
@@
(1) We introduce MobileGeo, a novel method built upon the precision-focused efficient design paradigm. This approach resolves the accuracy-efficiency dilemma in CVGL by concentrating model complexity during training, yielding a highly accurate inference model that is lightweight enough for deployment on mobile resource-constrained devices.
@@
(3) We propose the Multi-view Selection Refinement Module (MSRM) and provide a theoretical demonstration grounded in mutual information explaining how it enhances localization by optimally selecting and fusing multi-view information while minimizing feature matching overhead.
(4) We conduct extensive empirical evaluations on widely used benchmarks, including University-1652 and SUES-200, demonstrating that MobileGeo establishes a new state-of-the-art in both accuracy and efficiency. Additional deployment on edge devices further validate its real-time capabilities in the context of Mobile Edge Computing (MEC).
@@ II Related work
II-A Cross-view drone Geo-localization
Cross-View Geo-Localization involves matching a query image against a geo-referenced database from another view. Early methods relied on handcrafted descriptors like SIFT [16] or global scene descriptors like HoG [17], but showed limited robustness to the extreme cross-view domain gap. The paradigm shifted with deep metric learning, where Siamese or Triplet networks [18, 19] learn a shared, view-invariant embedding space. To explicitly address geometric inconsistencies, some works integrate geometric priors, notably using a polar transform to align satellite and ground-level perspectives [20, 21] before feature comparison. While ground-to-satellite matching [22, 23] sets the foundation, the unique oblique perspective of drone introduces a more severe and spatially non-uniform multi-view set of geometric challenges.
Drone-to-Satellite Geo-Localization. This task is particularly challenging due to the multi-view oblique and low-altitude perspective of drone, which creates significant domain gaps between platforms. Following the establishment of key benchmarks [1, 13], research has progressed from improving feature robustness with attention mechanisms [24] to the now-dominant Vision Transformer (ViT) architectures. ViTs like TransGeo [7] and FSRA [25] leverage self-attention to learn powerful global representations, setting high performance standards. More recently, the field has explored more efficient and powerful backbone architectures. For instance, several works have successfully employed ConvNeXt [26] to extract highly discriminative global features, achieving a strong balance of performance and efficiency [27]. This trend highlights a focus on backbone potency for feature extraction. However, most of contemporary methods [4, 10, 14] achieve higher accuracy by incorporating sophisticated auxiliary branches or modules, this paradigm is ill-suited for resource-constrained platforms. We follow a precision-focused efficient design paradigm to empower a simple backbone. By strengthening its feature extraction capabilities without adding any inference-time modules, our model achieves competitive precision while drastically reducing the computational burden, making it ideal for deployment on platforms like drone.
@@ II-B Multi-view Refinement
Traditional drone-based geo-localization relies on direct matching between individual query images and a reference database [28, 3, 8]. However, these approaches struggle with significant viewpoint discrepancies caused by factors such as occlusions from structures or vegetation, and the diverse perspectives captured by drone operating at varying altitudes and angles, making accurate location recognition increasingly challenging. Recent works have recognized the importance of leveraging multiple drone views to improve localization accuracy with approaches using 3D reconstruction [29, 30, 31] to represent scenes from multi-view observations and iteratively refining camera poses to align rendered views with satellite imagery. This multi-view fusion paradigm [32] demonstrates significant improvements in localization performance by exploiting the rich geometric and appearance information contained across different viewpoints.
@@
Mutual information (MI) has emerged as a principled criterion for selecting informative views in a wide variety of complex systems. In 3D reconstruction, MI has guided next-best-view selection effectively [33]. Similarly, in multi-view clustering, recent work [34] minimizes MI between common and view-specific representations to exploit inter-view complementary information while maximizing MI between refined and original representations to preserve principal information.
Existing drone-based geo-localization methods process all available views through expensive reconstruction-based methods [15, 29], missing the opportunity for principled information-theoretic view selection. However, despite their theoretical advantages, these multi-view approaches impose substantial computational burdens that severely limit their practical drone onboard applicability. In contrast to these complex approaches with substantial computational barriers, we propose an efficient MSRM that operates as a lightweight post-processing method. Rather than constructing expensive 3D representations or performing iterative rendering optimization, our method directly aggregates selected features from multiple drone viewpoints through Mutual Information (MI).
II-C Efficient On-Device Inference for Mobile Systems
The challenge of deploying CVGL models on mobile devices is a central theme in mobile computing research [35, 36]. Knowledge Distillation (KD) has become an essential technique in transferring knowledge from large, powerful models to efficient ones, particularly in environments with constrained resources, such as Mobile Edge Computing (MEC). Several approaches [37, 38] have explored this concept in the context of edge intelligence. For example, Xu et al. [39] proposed a heterogeneous federated learning framework, driven by multi-knowledge distillation, which facilitates efficient knowledge transfer between models with different architectures. In another study, Zhang et al. [40] introduced a bidirectional knowledge distillation method, enhancing the knowledge transfer process by simultaneously distilling knowledge from both the teacher and student models. Additionally, Luo et al. [41] proposed KeepEdge, a knowledge distillation-empowered edge intelligence framework for UAV delivery. Their work demonstrates the practical application of KD in improving edge intelligence for UAVs in MEC systems.
As mentioned above, the cross model distillation from a foundation model provides general-purpose guidance, but may not fully adapt to the target task’s specific constraints. Conversely, self-distillation [42] ensures internal consistency but lacks an external source of ”expert” knowledge. Our work is situated at the intersection of these two paradigms. We propose a Hierarchical Knowledge Distillation (HD-CVGL) paradigm, enhancing robustness and generalization without requiring any external modules during inference.
@@ III Proposed Method
In this section, we provide a comprehensive introduction to our proposed MobileGeo. As illustrated in Figure 3, we follow a precision-focused efficient design in two stages: training and inference. During the training stage, we propose Hierarchical Distillation for CVGL (HD-CVGL). Initially, we employ Fine-Grained Inverse Self-distillation (FISD) based on the internal stages of an efficient student model. Concurrently, to address the challenge of data imbalance, we introduce Uncertainty-Aware Prediction Alignment (UAPA), achieving a more robust and stable train. Subsequently, guided by a fine-tuned foundation model, we perform cross-distillation to obtain a student model with higher precision. In the inference stage, we propose a Multi-view Selection Refinement Module (MSRM) to better leverage multi-view information and further reduce the time cost of feature matching.
@@ III-A Hierarchical Distillation for CVGL
The second step of our hierarchical framework is Cross-Distillation Training (CDT), a process designed to transfer knowledge from a large foundation model (DINOv2-base [47]) teacher to a lightweight ConvNeXt-tiny [26] student, as illustrated in Figure 3 (a). This approach adheres to the MobileGeo paradigm by directly training the final student architecture without requiring any inference-time modifications. Critically, this teacher is not used off-the-shelf; it is first specialized for the target task through a parameter-efficient fine-tuning process on the University-1652 dataset, where only the final two Transformer blocks were made trainable. This approach creates an expert teacher that retains vast, general visual knowledge while acquiring high-level semantic understanding specific to our domain.
@@
III-B Multi-view Selection Refinement Module (MSRM)
@@
To address this challenge, as shown in Figure 5, we propose the MSRM, a post-processing technique that intelligently selects an optimal subset with . The key insight is that not all views contribute equally to geo-localization accuracy: some perspectives capture more distinctive features, while others may be less informative due to occlusions or viewing angles. We formulate this view selection problem within an information-theoretic framework, maximizing the mutual information between landmark identities while ensuring spatial diversity.
@@
Information-Theoretic View Importance Assessment. Given extracted multi-view features for a landmark, the MSRM quantifies each view’s information content through three complementary measures grounded in information theory. First, we approximate the marginal entropy of view features through log-variance:
@@
Computational Efficiency: By reducing the feature matching complexity from to , our MSRM achieves a significant speedup. is the number of drone views of one landmark, this corresponds to a reduction in computational cost, rendering the approach viable for real-time applications.
@@ IV-A Implementation Details
Evaluation Protocol. Following standard practice in cross-view geo-localization [1, 7], we evaluate both retrieval accuracy and computational efficiency across all models.
@@
Efficiency Metrics. We report three efficiency metrics: (1) Parameters: the total number of learnable parameters in millions; (2) FLOPs: floating-point operations computed for a single forward pass; (3) Processing Speed (FPS): in the feature extraction context, the number of drone and satellite images processed per second. Note that we set the batch size to 16 to accommodate the multi-view drone data processing.
@@
| LPN [2] (Wang et al. 2021) | 62.39 M | 65.39 G | 218 | 77.71 | 80.80 | 90.30 | 78.78 |
| FSRA [8] (Dai et al. 2021) | 53.16 M | 98.05 G | 100 | 85.50 | 87.53 | 89.73 | 84.94 |
| MCCG [28] (Shen et al. 2023) | 56.65 M | 51.04 G | 313 | 89.40 | 91.07 | 95.01 | 89.93 |
| MuSe-Net [53] (Wang et al. 2024) | 82.90 M | 42.37 G | 405 | 74.48 | 77.83 | 88.02 | 75.10 |
| SCPNet [54] (Gao et al. 2025) | - | - | - | 79.96 | 83.04 | 87.33 | 79.87 |
| TriSA [3] (Sun et al. 2024) | 51.13 M | 43.18 G | 275 | 90.08 | 91.56 | 96.01 | 90.12 |
| Safe-Net [55] (Lin et al. 2024) | 52.67 M | 24.58 G | 282 | 86.98 | 88.85 | 91.22 | 86.06 |
| SDPL [9] (Chen et al. 2024) | 42.56 M | 69.71 G | 519 | 90.16 | 91.64 | 93.58 | 89.45 |
| SRLN [56] (Lv et al. 2024) | 193.03 M | - | - | 92.70 | 93.77 | 95.14 | 91.97 |
| Sample4Geo [14] (Deuser et al. 2023) | 87.57 M | 90.24 G | 144 | 92.65 | 93.81 | 95.14 | 91.39 |
| CCR [11] (Du et al. 2024) | 156.57 M | 160.61 G | - | 92.54 | 93.78 | 95.15 | 91.80 |
| DAC [10] (Xia et al. 2024) | 96.50 M | 90.24 G | 128 | 94.67 | 95.50 | 96.43 | 93.79 |
| MEAN [4] (Chen et al. 2025) | 36.50 M | 26.18 G | 307 | 93.55 | 94.53 | 96.01 | 92.08 |
| MobileGeo (Ours) | 28.57 M | 4.45 G | 1022 | 93.87 | 94.83 | 95.72 | 92.57 |
| MobileGeo (Ours w/ Post-process) | 28.57 M | 4.45 G | 1022 | 97.15 | 97.50 | 95.58 | 96.27 |